# Adaptive and robust watermark against model extraction attack

### v1

@@ Abstract
Large language models have boosted Large Models as a Service (LMaaS) into a thriving business sector. But even model owners offering only API access while keeping model parameters and internal workings private, their Intellectual Property (IP) are still at risk of theft through model extraction attacks. To safeguard the IP of these models and mitigate unfair competition in the language model market, watermarking technology serves as an efficient post-hoc solution for identifying IP infringements. However, existing IP protection watermarking methods either explicitly alter the original output of the language model or implant watermark signals in the model logits. These methods forcefully distort the original distribution of the language model and impact the sampling process, leading to a decline in the quality of the generated text. The existing method also fails to achieve end-to-end adaptive watermark embedding and lack robustness verification in complex scenarios where watermark detection is subject to interference. To overcome these challenges, we propose PromptShield, a plug-and-play IP protection watermarking method to resist model extraction attacks without training additional modules. Leveraging the self-reminding properties inherent in large language models, we encapsulate the user’s query with a watermark self-generated instruction, nudging the LLMs to automatically generate watermark words in its output without compromising generation quality. Our method does not require access to the model’s internal logits and minimizes alterations to the model’s distribution using prompt-guided cues. Comprehensive experimental results consistently demonstrate the effectiveness, harmlessness, and robustness of our watermark. Moreover, Our watermark detection method remains robust and high detection sensitivity even when subjected to interference.
@@
Large Language Models, Model Extraction Attack, Watermarking
@@ I Introduction
In the current era of Language Model as a Service (LMaaS), safeguarding the intellectual property (IP) of large language models (LLMs) becomes a critical issue. The training of LLMs demands significant intellectual efforts, involving massive amounts of high-quality data, extensive computational resources, and human-elaborated training design. These investments imbue LLMs with substantial knowledge value, necessitating their inclusion under the IP rights of the legal owner.
@@
Despite the rapid growth of the LMaaS market, the development of comprehensive legal frameworks and regulations for IP protection lags behind [1, 2, 3]. The service process of LLMs is vulnerable to intellectual property infringement through model extraction attacks, also referred to as model stealing or imitation attacks. Model extraction attacks are akin to knowledge distillation [4], illustrated by the Alpaca case [5]. These attacks create a functionally comparable model in specific domains by distilling the victim model’s knowledge based on the queried victim model’s outputs [6, 7]. Such practices, especially when victim LLMs’ outputs are used for commercial model training, constitute a clear IP infringement. OpenAI’s policy [^1] Terms of use: https://openai.com/policies/terms-of-use, for instance, prohibits using ChatGPT’s outputs for competitive model training, underscoring the legal complexities and the need for stringent IP protection in AI development. Some top AI companies even have been caught stealing their competitors’ model outputs for training [8].
To avoid language model IP copyrights being offended, several real-time proactive methods for detecting attackers [9, 10, 11] and post-attack verification methods [12, 13, 14, 15, 16] for ownership authentication have been proposed. Watermarking, an information-hiding technique that embeds secret messages into carrier data, has been widely used for ownership protection [17, 18, 19, 20]. The purpose of watermarking is not to prevent IP stealing, as we cannot stop the attacker from imitating the actions of a normal user. Watermarks serve as a passive verification method, enabling the post-theft verification of IP ownership and complementing proactive defense strategies.
@@
Existing IP protection watermarking methods typically involve embedding special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership by designing heuristic watermark rules [15, 14, 16]. For example, He [15] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Such token replacement visibly alters the original language model’s prediction results at the token level. Zhao [16] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. This method also modifies the original distribution of the language model at the token probabilities level. Most existing IP protection watermarks are designed by modifying the distribution of language models through heuristic watermarking rules to obtain watermarking space. However, this inevitably distorts the original distribution of the language model and impacts the sampling process, resulting in the final sample deviating from optimal sampling and subsequently affecting the model’s performance in its downstream tasks. Also, these heuristic watermarking rules also lack sufficient flexibility and adaptability in real-world scenarios, struggling with unpredictable user queries that result in issues either outside the range of the watermark word list or with entropy too low to be effectively grouped.
Moreover, there is a lack of robust watermark detection in complex scenarios. Once attackers have obtained responses from the victim model in real-world scenarios, they can launch unpredictable and diverse attacks on the watermarked data. Attackers can also edit or replace watermarked text, which may nullify watermarks that rely on fixed synonym replacement rules. Discussions in content watermarking [21, 22] often highlight edit attacks at both the text token-level and sentence-level. Watermark robustness in scenarios of anti-model extraction attacks should also be resilient against potential dilution processes by attackers. Attackers might train imitation models using only a subset of query data or incorporate unwatermarked data. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermarks, or have only accounted for robustness against specific attacks. This oversight is evident both in the lack of robustness in the watermark design mechanism and the limited empirical robustness demonstrated in experiments.
To address these drawbacks, we propose PromptShield, a simple but effective watermarking method for language model IP protection that avoids forcefully distorting the language model distribution and has minimal impact on the original task quality. We encapsulate the user’s query into the language model’s self-watermarking system prompt, which automatically cues the model to embed watermarks. Due to their stable instruction understanding abilities, LLMs can seamlessly incorporate watermarks while formulating responses. We initially explore the capability of large language models to self-watermark, and to our knowledge, we are the first to empower language models to autonomously generate watermarks. Our findings confirm the viability of this auto-generated watermarking method, underscored by the language models’ capacity to learn and incorporate these watermarks effectively.
Our contributions include:
@@
  • We introduce a new plug-and-play adaptive watermarking method to combat model extraction attacks without additional training modules. Our method embeds watermarks automatically and directly during the inference stage without distorting language model distribution or disturbing the decoding, preserving the original performance of language models on downstream tasks.
  • We propose a robust watermark detection algorithm capable of effectively identifying watermarks even in realistic scenarios where the watermarks are subjected to interference. Remarkably, even if only 10 percent of watermarked data is utilized in training attackers’ models, our detector can confirm the unauthorized use of the protected model’s output.
@@
  • Through extensive experimentation—including scenarios with different watermark detection strategies, different epochs, and robustness tests, we have validated the effectiveness, learnability, harmlessness, and robustness of our watermarks.
@@ II-A1 Types of APIs
Model extraction attacks on victim language models vary based on the API service type. These attacks target label predictions, output embeddings, or output tokens.
@@ II-B1 Real-time defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [12, 13]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. As shown in Fig 12 in appendix VI-A, we categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks. We emphasize that the focus of our work is the IP protection for LLMs against model extraction attack, diverging from parameter watermarking [36, 37, 38, 39] to prevent the theft of the model parameters and content watermarking [21, 22, 40, 41] proving the ownership of model outputs. Specific introductions about the model parameter watermarks, content watermarks and dataset watermarks types, as well as the relationships between these watermarks, are presented in the appendix VI-A.
@@ II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use these watermarked model outputs to train their imitation models, these special signals can be detected in the output of the imitation models, allowing for effective IP protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflow of these two types of methods is demonstrated in the appendix VI-B.
White-box methods require knowledge of the internal details of the model to be protected and the ability to manipulate the text generation process. The representative white-box method is GINSEW [16], which groups and manipulates the conditional token probabilities of language model. It embeds specific frequency sine perturbations as watermark signals in the grouped probabilities, enabling the imitation models to learn this signal pattern during training. Watermark detection involves querying the imitation model for token probabilities, conducting matching filtering to detect grouped conditional probability outputs, and checking for the presence of embedded specific frequency signals. Similarly, the EmbMarker [42] uses a set of medium-frequency trigger words as watermark signals. EmbMarker selects a target embedding as the watermark and inserts it into the embeddings of texts containing trigger words as a backdoor. However, it may not be suitable for the current LLM market compared to the previous era of EaaS.
@@ III-A Preliminary
    Victim Model $M_{V}$: The object of the model extraction attacker’s attack, where the attacker can obtain the output from the victim’s model; it is also the object protected by our watermarking method, also known as the protected model.
@@
    Imitation Model $M_{I}$: An imitation model is a model trained with data from a victim model with the goal of becoming a replica of the victim model that is comparable in performance to the victim model in some way, which we also refer to as a thief model, mimic model.
@@ III-B The basic property of language model IP protection watermarks
As far as we know, existing research on watermarks against language model extraction attacks lacks a comprehensive description outlining essential watermark characteristics. We propose four fundamental properties crucial for effective watermarks in model extraction attack scenarios:
@@
III-C Adaptive Self-Watermarking Mechansim
Our goal is to devise a straightforward yet effective method for automatic watermark generation, avoiding manual intervention and extensive training processes like fine-tuning or Reinforcement Learning from Human Feedback (RLHF [44]). Given that existing LLMs possess an almost human-like capacity for understanding and adhering to instructions [45, 46, 47], our strategy leverages this strong instruction-following ability as a defensive mechanism in system mode. By utilizing the inherent proficiency of LLMs in token selection [48], we aim to guide them autonomously in generating watermarks to defend model extraction attacks. The watermark generation process resembles a person reminding themselves to incorporate a unique watermark known only to them during the spontaneous creation stage [49, 50].
Building on this insight, we introduce an additive system-mode self-watermarking method. We explore the capability of LLMs to autonomously and independently generate watermarks. Our method utilizes a system-generated prompt that instructs the LLM to embed an appropriate watermark based on its own assessment. We implement a system prompt that integrates the user query in the following manner:
[SYSTEM PROMPT] Use some special words as watermarks in the generated texts. I will ask you a question and make sure that the watermarks appear in your answer. Here is the question: [USER QUERY]
This system-mode self-watermarking approach is readily applicable across various LLMs, serving as an efficient identifier against model extraction without compromising the original output quality of the LLMs. In our experiments with several similar prompts, we observed comparable watermark generation effects as shown in the Table I. We find that due to the language model’s powerful prompt understanding capabilities, our automatic watermarking method is insensitive to the specific prompt used. While optimizing the automatic prompts might enhance performance in other downstream tasks, we don’t consider prompt searching as the focus of our work. Clearly, articulated prompts with similar intentions can also achieve similar effects in generating watermark words.
@@ III-C Adaptive Self-Watermarking Mechansim
TABLE I: System instructions
III-D Robust Model Watermark Detection
When model owners suspect a model theft, Firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. It is a common assumption in model stealing verification followed by [16] , also there is evidence that OpenAI [8] actually has the query history of every user’s query and answers that OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model. When an individual is suspected of using query data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{ori}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. However, the attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
@@ III-D Robust Model Watermark Detection
For the attacker, both ratios $\rho$ and $\eta$, are known and controllable. However, the watermark detector has no knowledge of these two ratios. Formally, the attacker has access to $\{Q,A,D,Q^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,Y_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $Y_{S}$ denotes the outputs from the imitation model using the query sets $Q$.
@@
Prior Rapid Watermark Verification: This method rapidly judges based on the outputs $W_{S}$ of the suspect model $M_{S}$ to $Q$. It enables quick and effective watermark identification without relying on machine learning modules.
Detailed Contrastive Watermark Verification: With knowledge of the suspect model’s base model information, we can further compare the output differences between the suspect model and the base model using the Kolmogorov-Smirnov (KS) test [51].
In prior rapid verification, we only need the outputs $Y_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output.
@@
where $WM_{i}$ is the watermark word set for each query $q_{i}$ and $\mathbbm{1}_{y_{i}}(w)$ is the indicator function for the presence of $w$ in $y_{i}$. $l(\cdot)$ is the function to count tokens length. In other words, we calculate the total length of all unique watermark words found in the suspect model’s generated text, normalized by the length of the generated text. The average Sentence Watermark Score (ASWS) is calculated by ${\sum_{i=1}^{n}{S_{i}}}/{n}$.
@@
(a) Main Results from Using GPT2-Large as the Base Model on HC3
(b) Main Results from Using LLaMA2 as the Base Model on HC3
(c) Main Results from Using Mistral as the Base Model on HC3
(d) Main Results from Using GPT2-Large as the Base Model on WILD
(e) Main Results from Using LLaMA2 as the Base Model on WILD
(f) Main Results from Using Mistral as the Base Model on WILD
Figure 3: Main Experiment Results: We trained imitation models ($M_{I}$) using watermarked data based on three different foundational models, demonstrating the average sentence watermark score from all texts generated by the imitation models and their question-answering performance. Compared to outputs from the original foundational models ($M_{O}$) and models trained with normal data ($M_{L}$), the watermark scores of the imitation models were significantly higher, while their question-answering performance remained comparable to that of normal models trained without watermarked data ($M_{L}$).
In the detailed contrastive watermark verification, we employ two-sample KS test [51], which is inspired by the detection method in Membership Inference Attacks [53]. The KS statistic is the maximum difference between the cumulative distribution functions (CDFs) of the two samples. Given the distribution $D_{s}$ of Sentence Watermark Scores from the outputs of suspect model $M_{s}$, the distribution $D_{L}$ from the model trained with unwatermarked data $M_{L}$, the distribution $D_{O}$ from the origin base language model $M_{O}$, we intuitively think the Sentence Watermark Score distribution of the imitation model $M_{I}$ should have apparent difference from $M_{L}$ and $M_{O}$. The KS test evaluates whether these distributions exhibit significant differences, with an emphasis on their cumulative properties. A low p-value from the KS test indicates that the distributions are significantly different, while a high p-value indicates that the distributions are similar and consistent with each other. the KS statistic defined for $M_{L}$ ,$KS_{L}$ is defined as:

### v2

# ModelShield: Adaptive and Robust Watermark against Model Extraction Attack

@@ Abstract
Large language models (LLMs) demonstrate general intelligence across a variety of machine learning tasks, thereby enhancing the commercial value of their intellectual property (IP). To protect this IP, model owners typically allow user access only in a black-box manner, however, adversaries can still utilize model extraction attacks to steal the model intelligence encoded in model generation. Watermarking technology offers a promising solution for defending against such attacks by embedding unique identifiers into the model-generated content. However, existing watermarking methods often compromise the quality of generated content due to heuristic alterations and lack robust mechanisms to counteract adversarial strategies, thus limiting their practicality in real-world scenarios. In this paper, we introduce an adaptive and robust watermarking method (named ModelShield) to protect the IP of LLMs. Our method incorporates a self-watermarking mechanism that allows LLMs to autonomously insert watermarks into their generated content to avoid the degradation of model content. We also propose a robust watermark detection mechanism capable of effectively identifying watermark signals under the interference of varying adversarial strategies. Besides, ModelShield is a plug-and-play method that does not require additional model training, enhancing its applicability in LLM deployments. Extensive evaluations on two real-world datasets and three LLMs demonstrate that our method surpasses existing methods in terms of defense effectiveness and robustness while significantly reducing the degradation of watermarking on the model-generated content.
@@
Large Language Models, Model Extraction Attack, Text Watermarking, Model IP Protection
@@ I Introduction
In the current era of Language Model as a Service (LMaaS), safeguarding the intellectual property (IP) of large language models (LLMs) becomes a critical issue. The training of LLMs demands significant intellectual efforts, involving massive amounts of high-quality data, extensive computational resources, and human-elaborated training design. These investments imbue LLMs with substantial knowledge value, necessitating their inclusion under the IP rights of the legal owner. Unfortunately, the service process of LLMs is vulnerable to intellectual property infringement through model extraction attacks, also referred to as model stealing or imitation attacks. Model extraction attacks are akin to knowledge distillation [1], illustrated by the Alpaca case [2]. These attacks create a functionally comparable model in specific domains by distilling the victim model’s knowledge based on the queried victim model’s outputs [3]. Such practices, especially when victim LLMs’ outputs are used for commercial model training, constitute a clear IP infringement. OpenAI’s policy [^1] Terms of use: https://openai.com/policies/terms-of-use, for instance, prohibits using ChatGPT’s outputs for competitive model training, underscoring the legal complexities and the need for stringent IP protection in AI development.
@@
Watermarking technique [4, 5, 6] can embed secret messages into target data, which has been widely used for protecting the intellectual property of the AI models. Existing watermarking methods for IP protection typically follows a similar framework, which utilizes heuristic watermark strategies to embed special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership [7, 8, 9]. For example, He [7] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Zhao [9] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. In conclusion, most existing watermark methods for IP protection embed secret message by modifying the distribution of language models through heuristic watermarking rules. However, this inevitably distorts the original distribution of the language model and impacts the normal generation process, resulting in the final generation results deviating from the optimal one and subsequently affecting the model’s performance in its downstream tasks.
Moreover, the attackers can employ diverse adversarial techniques to attack the content generated by the victim model to degrade the effectiveness of the watermark methods. In one hand, attackers are able to modify the watermarked texts, which can reduce the sensitivity of the embedded watermark signals, especially for the watermark methods that rely on fixed synonym replacement rules [7, 8]. In another hand, attackers may also select a part of the watermarked content, and mix them with some clean data to train the imitation model, which can further reduce the impact of the watermarked data on behavior of the imitation model. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermark strategies and the corresponding detection mechanisms, or have only accounted for robustness against specific attacks.
To address these drawbacks, we propose ModelShield, a simple yet effective watermarking method for language model IP protection, which can embed watermarks into model generations without degrading the accuracy of generated content, meanwhile effectively detecting the watermark embedded in the imitation model under the interference of adversarial techniques. Specifically, to avoid distorting the generation distribution of language model via manually-designed rules, we combine the user’s query with a self-watermarking prompt to instruct the victim LLM, which can empower the LLM to adaptively embed watermarks into the generated content. Besides, we further propose a robust watermark detection method to model the strength of watermark signal embedded in content generated by the imitation model and determine the IP infringement. This method also introduces a robust mechanism by filtering the disturbance samples based on the strength of watermark signal encoded in them, to defend against the potential adversarial attacks, like editing or dilution attacks. We conduct extensive experiments on two benchmark QA datasets [10, 11], using ChatGPT as the victim model and GPT-2 large [12], LLama2 [13], and Mistral [14] as the base models for imitation. Results show that our proposed ModelShiled method can adaptively and effectively embed the watermark signals into the generation content, meanwhile reducing the adverse effects of watermarks on the quality of the generated content by 29.08%. Results also demonstrate ModelShied maintains the effectiveness of the watermark while preserving the performance of the language model when adversarial techniques are incorporated to attack the watermarked content.
The contribution of this paper are summarized into three-fold:
@@
  • We introduce a new plug-and-play adaptive watermarking method to combat model extraction attacks without additional training modules. Our method can adaptively embed watermarks messages based on the generated content without manually distorting language model distribution, which can preserve the original performance of language models on downstream tasks.
  • We propose a robust watermark detection algorithm that exhibits high sensitivity to watermark signals embedded within victim models, even when trained on a small subset of user query data (e.g., 400 queries). Our method also maintains its effectiveness despite the presence of various adversarial techniques.
@@
  • Extensive experiments, including scenarios with various watermark detection strategies, different epochs, and robustness tests on two benchmark datasets and three foundational models, demonstrate the effectiveness, harmlessness, and robustness of our ModelShield method.
@@ II-A1 Types of APIs
Model extraction attacks on language models vary based on the API service type. These attacks target label predictions, output embeddings, or output tokens.
@@ II-B1 Real-time defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [31, 32]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. As shown in Fig 12 in supplementary VI-A, we categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks. We emphasize that the focus of our work is the IP protection for LLMs against model extraction attack, diverging from parameter watermarking [33, 34, 35, 36] to prevent the theft of the model parameters and content watermarking [37, 38, 39, 40] proving the ownership of model outputs. Specific introductions about the model parameter watermarks, content watermarks and dataset watermarks, as well as the relationships between these watermarks, are presented in the supplementary VI-A.
@@ II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use these watermarked model outputs to train their imitation models, these special signals can be detected in the output of the imitation models, allowing for effective IP protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflows of these two types are demonstrated in the supplementary VI-B.
White-box methods require knowledge of the internal details of the model to be protected and the ability to manipulate the text generation process. The representative white-box method is GINSEW [9], which groups and manipulates the conditional token probabilities of language model. It embeds specific frequency sine perturbations as watermark signals in the grouped probabilities, enabling the imitation models to learn this signal pattern during training. Watermark detection involves querying the imitation model for token probabilities, conducting matching filtering to detect grouped conditional probability outputs, and checking for the presence of embedded specific frequency signals. Similarly, the EmbMarker [41] uses a set of medium-frequency trigger words as watermark signals. EmbMarker selects a target embedding as the watermark and inserts it into the embeddings of texts containing trigger words as a backdoor. However, it is not suitable for the current LLM market compared to the previous era of EaaS.
@@ III-A Preliminary
    Victim Model $M_{V}$: The target of the model extraction attack, where the attacker can obtain the output from the victim’s model; it is also the object protected by our watermarking method, also known as the protected model.
@@
    Imitation Model $M_{I}$: An imitation model is a model trained with data from a victim model with the goal of becoming a replica of the victim model that is comparable in performance to the victim model in some way, which is also referred to as a thief model, mimic model.
@@ III-B The basic property of language model IP protection watermarks
We propose four fundamental properties crucial for effective watermarks in model extraction attack scenarios:
@@
III-C Adaptive Self-Watermarking Mechanism
Our goal is to devise a straightforward yet effective method for automatic watermark generation, avoiding manual intervention and extensive training processes like pre-training or fine-tuning. Given that existing LLMs possess an almost human-like capacity for understanding and adhering to instructions [43, 44], our strategy leverages this strong instruction-following ability as a defensive mechanism in system mode. By utilizing the inherent proficiency of LLMs in token selection [45], we aim to guide them autonomously in generating watermarks to defend model extraction attacks. The watermark generation process resembles a person reminding themselves to incorporate a unique watermark known only to them during the spontaneous creation stage [46, 47].
Building on this insight, we introduce an additive system-mode self-watermarking method. We explore the capability of LLMs to autonomously and independently generate watermarks. Our method utilizes a system-generated prompt that instructs the LLM to embed an appropriate watermark based on its own assessment. For example, we implement a system prompt that integrates the user query in the following manner:
[USER QUERY]: Write a email to your boss explaining why you should have a pay raise. [SYSTEM PROMPT]: Use specific strategies and include some special words as watermarks in your generated text. Ensure these watermarks are present in your answer.
This system-mode self-watermarking approach is readily applicable across various LLMs, serving as an efficient identifier against model extraction without compromising the original output quality of the LLMs. In our experiments with several similar prompts, we observed comparable watermark generation effects as shown in the Table I. We find that due to the language model’s powerful prompt understanding capabilities, our automatic watermarking method is insensitive to the specific prompt used. While optimizing the automatic prompts might enhance performance in other downstream tasks, we don’t consider prompt searching as the focus of our work. Clearly, articulated prompts with similar intentions can also achieve similar effects in generating watermark words. Afterward, the model owner can format the generated text, returning only relevant answers to the user and recording the watermark words from the query.
@@ III-C Adaptive Self-Watermarking Mechanism
TABLE I: System instructions for enabling the victim model to generate watermarks automatically.
III-D Robust IP Infringement Detection
When model owners suspect a model theft, firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. It is a common assumption in model stealing verification followed by [9] , also there is evidence that OpenAI [48] actually has the query history of every user’s query and answers that OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model.
When an individual is suspected of using query data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{O}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. The attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
@@ III-D Robust IP Infringement Detection
For the attacker, both ratios $\rho$ and $\eta$, are known and controllable. However, the watermark detector has no knowledge of these two ratios.
Formally, the attacker has access to $\{Q,A,D,Q^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,Y_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $Y_{S}$ denotes the outputs from the imitation model using the query sets $Q$.
@@
Prior Rapid Watermark Verification:
This method rapidly judges based on the outputs $W_{S}$ of the suspect model $M_{S}$ to $Q$. It enables quick and effective watermark identification without relying on machine learning modules.
Detailed Contrastive Watermark Verification:
With knowledge of the suspect model’s base model information, we can further compare the output differences between the suspect model and the base model using the Kolmogorov-Smirnov (KS) test [49]. Furthermore, watermark detectors can train the base model using unwatermarked data on the known base model, allowing for a more detailed comparison with the suspect model.
In prior rapid verification, we only need the outputs $Y_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output $W_{W}$.
@@
where $WM_{i}$ is the watermark word set for query $q_{i}$ and $\mathbbm{1}_{y_{i}}(w)$ is the indicator function for the presence of $w$ in $y_{i}$. $l(\cdot)$ is the function to count tokens length. In other words, we calculate the total length of all unique watermark words found in the suspect model’s generated text, normalized by the length of the generated text. The Average Sentence Watermark Score (ASWS) is calculated by ${\sum_{i=1}^{n}{S_{i}}}/{n}$.
@@
(a) Using Gpt2-Large as the base model on HC3
(b) Using Llama2 as the base model on HC3
(c) Using Mistral as the base model on HC3
(d) Using Gpt2-Large as the base model on WILD
(e) Using Llama2 as the base model on WILD
(f) Using Mistral as the base model on WILD
Figure 3: Main Experiment Results: We trained imitation models ($M_{I}$) using 4000 watermarked data based on three different foundational models, repeating the process ten times. The results demonstrate the average sentence watermark score from all texts generated by the imitation models, as well as their question-answering performance. Compared to outputs from the original foundational models ($M_{O}$) and models trained with normal data ($M_{L}$), the watermark scores of the imitation models were significantly higher, while their question-answering performance remained comparable to that of normal models trained without watermarked data ($M_{L}$).
In the detailed contrastive watermark verification, we employ two-sample KS test [49], which is inspired by the detection method in Membership Inference Attacks [51]. The KS statistic is the maximum difference between the cumulative distribution functions (CDFs) of the two samples. Given the distribution $D_{S}$ of Sentence Watermark Scores from the outputs of suspect model $M_{S}$, the distribution $D_{L}$ from the model trained with unwatermarked data $M_{L}$, the distribution $D_{O}$ from the origin base language model $M_{O}$, we intuitively think the Sentence Watermark Score distribution of the imitation model $M_{I}$ should have apparent difference from $M_{L}$ or $M_{O}$. The KS test evaluates whether these distributions exhibit significant differences, with an emphasis on their cumulative properties. A low p-value from the KS test indicates that the distributions are significantly different, while a high p-value indicates that the distributions are similar and consistent with each other. the KS statistic defined for $M_{L}$ ,$KS_{L}$ is defined as: